OpenStudy (anonymous):

translate into algebraic expression the product of three and a number increased by five

OpenStudy (anonymous):

Let \(n\) represent our number. How would we write three times that number (\(3 \times \))? How would we write that increased by five (\(+5\))?

OpenStudy (e.cociuba):

x=3+x*5

OpenStudy (e.cociuba):

x equals three, plus x, times 5

OpenStudy (anonymous):

@e.cociuba, that's not quite correct. You should have an expression, not an equation.

OpenStudy (amistre64):

the lack of commas makes this a bit ambiguous to me

OpenStudy (anonymous):

so it will be n=3+x

OpenStudy (e.cociuba):

k so then u help us all out here. im a bit confussed!

OpenStudy (amistre64):

no

OpenStudy (amistre64):

i see 2 ways of interpreting this .... which makes it ambiguous to me the product of: (three and a number) increased by five the product of: three and (a number increased by five)

OpenStudy (amistre64):

the first option is my gut instinct tho

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

how would you write the product of three and a number?

OpenStudy (anonymous):

times 3

OpenStudy (amistre64):

yes; but lets use 3*x just to be safe :) then just add five to it

OpenStudy (anonymous):

3*x+5

OpenStudy (amistre64):

good

OpenStudy (e.cociuba):

thats the same thing i did, i just did 3+x*5 (u just switched the 3 and 5)

OpenStudy (e.cociuba):

rite?

OpenStudy (amistre64):

your math sentence does not reflect the english wording.

OpenStudy (anonymous):

can u help me with an another

OpenStudy (anonymous):

fifteen less than twice a number

OpenStudy (amistre64):

yours represents: what is the sum of 3, and the product of 5 and a number

OpenStudy (amistre64):

how would you write: twice a number

OpenStudy (e.cociuba):

Oh well I’m sorry for confusing u guys with my English wording.

OpenStudy (anonymous):

i am not sure

OpenStudy (amistre64):

twice also can be read as "2 times" so (2 times a number) can be written as?

OpenStudy (anonymous):

n*2

OpenStudy (amistre64):

good; and when we less that by fifteen .... which means we subtract 15 from it we get?

OpenStudy (anonymous):

n*2-15

OpenStudy (amistre64):

correct. or another way that you can write it is: 2*x - 15

OpenStudy (amistre64):

good luck ;)